The wires in a household lamp cord are typically 3.5 mm apart center to center and carry equal currents in opposite directions.If the cord carries current to a 100−W light bulb connected across a 120−V potential difference, what force per meter does each wire of the cord exert on the other? (Model the lamp cord as a very long straight wire.)
Answer #1

We know that:
P = V*I
=> I = P / V = 100W / 120 = 833mA
So, FOrce per unit length is given by:
F / L = (meu(o) / 2pi) (I1I2 / d)
F / L = (4pi x 10^-7 T / 2pi) * ( 833mA * 833mA / 3.5mm) = 3.965 * 10^-5 N/meter